the green line in the atomic spectrum of thallium has a wavelength of 535 nm


1
Expert's answer
2022-01-05T13:44:02-0500


Question:

The green line in the atomic spectrum of the thallium atom has a wavelength of 535 nm. Calculate the energy, in joules, of the photon with this wavelength. (c = 3 * 10^8 m/s) & (h = 6.626 * 10^-34 J * s)



ANSWER


"E=hf"


But "F=\frac{c}{\lambda}"


"E=h\frac{c}{\lambda}"


"=\frac{6.626×10^{-34}×3×10^8}{535×10^{-9}}"


"=3.176\u00d710^{-19}J"